OpenStudy (anonymous):

The length of a rectangle is 2 inches more than the width. The perimeter is 40 inches. Find the length and width.

OpenStudy (anonymous):

l=2+w P=2(l+w)=40 s solve for l an w

OpenStudy (anonymous):

how? i dont understand

OpenStudy (anonymous):

perimeter=2(l+w) length=2+w perimeter=2(2+w+w) =2(2+2w) =4+4w=40 4w=36 w=36/4 w=9 l=2+w l=11

OpenStudy (anonymous):

The length of a rectangle is 7 inches more than three times the width. The perimeter is 94 inches. Find the length and width.

OpenStudy (anonymous):

l=2+w...........................1 2l+2w=40.....................2 plug in l from equation 1 into equation 2

OpenStudy (anonymous):

3h by 4w =5

OpenStudy (anonymous):

l=7+3w 2l+2w=94

OpenStudy (anonymous):

OK. let's draw a rectangle... We have length=7+3w and width=w. So I just labeled the length to contain a variable 'w' so we can make an equation. The definition of a perimeter is 2 times the length and width to make P= 2L + 2W. I substitute L or length= 7+3w since the question gave us "7 inches more than three times the width", which basically translates to that. So the equation becomes P= 2(7+3w) + 2w = 94. If you solve this equation for w, you'll get 8w=80, which makes w=10. We found the width! now we want to substitute w to our original equation length which was 7+3w. So 7+3(10)=37. So the answer is 37 for length and 10 for the width or whatever you'd like to call what is length of width. Hope this helps!

OpenStudy (anonymous):

oops, hang on. I think I made a mistake. lol. I forgot to factor in the perimeter. Hang on guys.....

OpenStudy (anonymous):

11+11+9+9=40

OpenStudy (anonymous):

Oh wait, there are two questions. I looked at the second one. The original question I did not answer. You can always put in the numbers back into the equations to get the total perimeter to get a total of 94 inches, which is the perimeter given in the second question.

OpenStudy (anonymous):



OpenStudy (raden_zaikaria):

understand sasha22 ....

OpenStudy (raden_zaikaria):

The perimeter is the total distance around the outside of a 2D shape

OpenStudy (anonymous):

11L x2+9Wx2=40

OpenStudy (anonymous):

the 3 4 5 is a trangent for another math question, sorry.